OpenStudy (anonymous):

Find the length of the radius of the given circle. (x - 5)2 + (y + 3)2 = 25

OpenStudy (radar):

An understanding of the circle equation is required to solve this problem.

OpenStudy (anonymous):

I personally need more of a visual. But I don't think you can put one on here, can you?

OpenStudy (radar):

The value to the right of the equal sign, is all that is needed to solve what is asked. 25 is what squared?

OpenStudy (anonymous):

625, right?

OpenStudy (radar):

However a complete understanding is needed to locate the center. (whiich was not asked)

OpenStudy (anonymous):

these are the chooses they give me (5, 3) (5, -3) (-5, 3)

OpenStudy (radar):

x^2 = 25 what is x?

OpenStudy (anonymous):

now I'm confused. I never could figure out what the ^ sign meant.

OpenStudy (radar):

625 is 25^2, but my wording was not specific, I should of said "what squared is equal to 25?"

OpenStudy (radar):

It indicated to what power the variable is raised. x^2 is x squared or x to the 2nd power, while x^3, means x is cubed or raised to the 3rd power.

OpenStudy (anonymous):

oooh okay. so ^ means the number is squared?

OpenStudy (anonymous):

(x - 5)^2 + (y + 3^)2 = 25

OpenStudy (radar):

Yes, if a 2 is next to the ^, x^2 means x is squared.

OpenStudy (anonymous):

oh okay. lol now you're helping both of us because I could never figure that out.

OpenStudy (radar):

The answers contain more information than was asked, The value 5 is the distance of the radius

OpenStudy (radar):

The value 3 is the offset of the y axis. Now revealing the complete formula (equation) for a circle: \[(x-h)^{2} + (y-k)^{2}= r ^{2}\]

OpenStudy (radar):

h is x offset, k is the y offset, by offset I mean value of the displacement from the origin.

OpenStudy (radar):

So the best answer is 5, -3, however I feel there is a typo, as the y offset was not asked for.

OpenStudy (anonymous):

they were supposed to ask for it?

OpenStudy (anonymous):

isn't the y offset 5?

OpenStudy (radar):

I don't know, it is difficult to speculate on the instructor's intention.

OpenStudy (radar):

The x offset is 5, so the origin is at (5,-3) My guess they are asking for the location of the center of the circle and not the length of the radius..........probably a typo.

OpenStudy (anonymous):

hmm okay.

OpenStudy (radar):

@adrian1220 Please review that problem, and determine what are they asking for.